OpenStudy (anonymous):

-x -7y=-34 x is -29 what's why explain using steps

OpenStudy (anonymous):

what's y is what that should read

OpenStudy (anonymous):

-x-7y=-34 substitute -(-29)-7y=-34 multiply 29-7y=-34 subtract -7y=-63 divide by -7 y=9

OpenStudy (anonymous):

1 more question

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

8x-2y=10 x is -1

OpenStudy (anonymous):

@mathman806

OpenStudy (anonymous):

a few more after that

OpenStudy (anonymous):

8x-2y=10 x=-1 substitute 8*-1-2y=10 multiply -8-2y=10 add 8 -2y=18 divide by -2 y=-9

OpenStudy (anonymous):

be back tomorrow